mathslover (mathslover):

Very urgent if 5th march 1995 is tuesday what was the day on 22nd october 1994

mathslover (mathslover):

@dpaInc

mathslover (mathslover):

@.Sam.

OpenStudy (anonymous):

the Rain Man can do this problem.... but he's watching Jeopardy... :(

mathslover (mathslover):

@radar

mathslover (mathslover):

@UnkleRhaukus

mathslover (mathslover):

unke rhaukus please help

OpenStudy (anonymous):

Can you wait I am working on it...

mathslover (mathslover):

ok sorry

OpenStudy (unklerhaukus):

do you want a method or jus the solution?

OpenStudy (anonymous):

I have a method but that is too long to explain...

mathslover (mathslover):

method also

OpenStudy (unklerhaukus):

i thought that the 5th march 1995 was a saturday

mathslover (mathslover):

no... it is not that of a real calender ... it is mental ability

OpenStudy (anonymous):

Is it Monday???

OpenStudy (anonymous):

I think the answer is Monday.... Is it @mathslover ??

mathslover (mathslover):

very right @waterineyes

mathslover (mathslover):

please check my soln : 5th march 1994 will be monday right?

mathslover (mathslover):

and then 26 + 30 + 31+30+31+31+30+22 = 231/7 = 0 odd days hence monday + 0 = monday hence 22nd october 1994

OpenStudy (unklerhaukus):

5th march 1995 is tuesday jan+feb+5 31+28+5 64=9x7+1 days past new-years eve so new years eve was a monday 22nd october 1994 31-22 +nov+dec 9+30+31 70 =7x10 so monday

mathslover (mathslover):

the year next to 2003 that has the same calander ? if it was sunday on 17 november 1962 what will be the day on 22nd november 1964 karan was born on saturday 22nd march 1982 on what day of the week was he was 14 year 7 month 8 days ....

mathslover (mathslover):

i have these more questions please help

OpenStudy (anonymous):

After 2003, 2014 will have the same calendar...

mathslover (mathslover):

how ?

OpenStudy (anonymous):

Just check that, in 2003 it is Wednesday on 1st January... After it only 2014 will have Wednesday on 1st January.. But remember, If 2003 is not a leap year then 2014 must not be the leap year and 2014 is not a leap year...

mathslover (mathslover):

hmn can u tell me that can i do this my this method : 2003 -1 2004 - 2 2005 - 1 2006 - 1 2007 - 1 2008 - 2 2009 -1 2010-1 2011-1 2012-2 2013-1 2014-1 total we get 15 odd days that is not that of 2003 ...

OpenStudy (anonymous):

Why are you including 2014 in it??

mathslover (mathslover):

why not include 2014 ?

OpenStudy (anonymous):

Is it Monday in 3rd question??

mathslover (mathslover):

i dont know there answers except that of 1 .... can u explain please

OpenStudy (anonymous):

Take a minute, If 1st Of Jan on 2013 is Wednesday, then what is the day on 1st Of Jan in 2014. Can you solve this question by using that ODD DAYS method just calculate and then decide how many odd days will you take???

mathslover (mathslover):

right .... i got it now ....it wednesday will be on 1st of jan 2014...

OpenStudy (anonymous):

Sorry there will come 2003 in place of 2013...

mathslover (mathslover):

i had just find the answer of 2nd ..... is that sunday ...

mathslover (mathslover):

np ... is my 2nd one correct ?

mathslover (mathslover):

22nd nov.1964 is sunday

mathslover (mathslover):

17 nov. 1964 will be tuesday and hence 22nd nov. 1964 will be sunday

OpenStudy (anonymous):

I will solve it then I will tell you.. Now mummy is calling me for a task.. I just come back in a few seconds...

mathslover (mathslover):

please come back soon i have to solve them urgently

mathslover (mathslover):

@UnkleRhaukus

OpenStudy (unklerhaukus):

hey

mathslover (mathslover):

can u help me for that 2 questions :if it was sunday on 17 november 1962 what will be the day on 22nd november 1964 karan was born on saturday 22nd march 1982 on what day of the week was he was 14 year 7 month 8 days ....

OpenStudy (unklerhaukus):

these problems are hurting my mind at this time of day

mathslover (mathslover):

ouch :(

mathslover (mathslover):

@waterineyes

OpenStudy (anonymous):

Ya I came back...Just wait I am solving at my level speed...

mathslover (mathslover):

k ..

OpenStudy (anonymous):

No, My answer came was Monday on 22nd November 1964 because 1964 is a leap year also...

mathslover (mathslover):

how ?

mathslover (mathslover):

@waterineyes

OpenStudy (anonymous):

Days between 17 Nov 1962 to 31 December 1962 = 13 + 31 = 44.. Odd days in 1963 is 1.. And number of days in 1964 is 327.. (44 + 327) = 371/7 gives 0 as Odd days... So total number of odd days is 1.. Therefore, I think Monday is the answer...

mathslover (mathslover):

k for 3rd one ?

mathslover (mathslover):

u there ?

mathslover (mathslover):

@waterineyes

OpenStudy (aravindg):

am i late ?

mathslover (mathslover):

can u answer me 3rd one @AravindG or 2nd and 3rd

mathslover (mathslover):

@waterineyes

OpenStudy (anonymous):

In third question, I think Saturday is the answer...